OpenStudy (anonymous):

Solve the following equations for all solutions of x 2cos(3x-1)=0 So do I just want to start out multiplying and getting 6cosx-2cos=0?

OpenStudy (anonymous):

not if it really is \[2\cos(3x-1)=0\][

OpenStudy (anonymous):

cosine is a function. think if it as \(f(x)=\cos(x)\) so you have \(f(3x-1)=\cos(3x-1)\)

OpenStudy (anonymous):

So you subtracted one of the cos(3x-1) from the other?

OpenStudy (lgbasallote):

you actually divide both sides by 2

OpenStudy (lgbasallote):

think of cos (3x - 1) as a variable y so you have 2y = 0 to get y alone, you divide both sides by 2 make sense?

OpenStudy (anonymous):

Yeah...so, do I end up with cos(3x-1)=0

OpenStudy (anonymous):

Think of this as cosA=0 and solve for A

OpenStudy (anonymous):

Ugg...this class is terrible!

OpenStudy (anonymous):

If cos A = 0 then A is 90 degees, 270 degrees and so on or pi/2 or 3pi/2 etc

OpenStudy (anonymous):

I really appreciate the help. Thanks everyone!

OpenStudy (anonymous):

Hope you do okay on this topic. Cheers for the medal

OpenStudy (anonymous):

Yeah...I've just been working n these all day and my mind is mush. Thanks again.